OpenStudy (anonymous):

Johnson and Johnson purchased a house for $450000. The had a 35% down payment and negotiated a four year mortgage at 4:35% amortized over 20 years. (a) Calculate their monthly payments. (b) After the four year term was completed, they renegotiated a new term for six years at 2:15% amortized over 15 years. Determine their new monthly payments.

OpenStudy (anonymous):

Step 1. 450000*.35 Step 2. 450000-157500 = 292500

OpenStudy (anonymous):

Ummm. And I don't understand how the rest is phrased...

OpenStudy (anonymous):

wow your a hero!

OpenStudy (anonymous):

can you please answer the whole thing lol ur a life saver honestly!

OpenStudy (anonymous):

Something with that equation. :[

OpenStudy (anonymous):

I don't understand how the rest is phrased let me read it again

OpenStudy (anonymous):

I don't understand the second part....what does 4:35% mean?

OpenStudy (anonymous):

is it 2329.4?

OpenStudy (anonymous):

for b?

OpenStudy (anonymous):

@_@ I'm confused. for part a. Was it only for a period of 4 years?

OpenStudy (anonymous):

wait nvm. over 20 years...but I don't understand that ratio thing

OpenStudy (anonymous):

hello? can you reply please

OpenStudy (anonymous):

Could you answer my question first lol. I really don't understand how your question is phrased

OpenStudy (anonymous):

yes now i can see my computer is messed up sorry

OpenStudy (anonymous):

Johnson and Johnson purchased a house for $450000. The had a 35% down payment and negotiated a four year mortgage at 4:35% amortized over 20 years. (a) Calculate their monthly payments. (b) After the four year term was completed, they renegotiated a new term for six years at 2:15% amortized over 15 years. Determine their new monthly payments.

OpenStudy (anonymous):

@_@ a four year mortgage over 20 years...how does that make sense :[

OpenStudy (anonymous):

i know right

OpenStudy (anonymous):

First you must define some variables to make it easier to set up: P = principal, the initial amount of the loan I = the annual interest rate (from 1 to 100 percent) L = length, the length (in years) of the loan, or at least the length over which the loan is amortized. The following assumes a typical conventional loan where the interest is compounded monthly. First I will define two more variables to make the calculations easier: J = monthly interest in decimal form = I / (12 x 100) N = number of months over which loan is amortized = L x 12 Okay now for the big monthly payment (M) formula, it is: J M = P x ------------------------ 1 - ( 1 + J ) ^ -N I found this on some website: http://www.hughchou.org/calc/formula.html

OpenStudy (anonymous):

but I don't know what that 4 year term means...

OpenStudy (anonymous):

that is throwing me off